# Transfer Learning Between U.S. Presidential Elections: How Should We Learn From A 2020 Ad Campaign To Inform 2024 Ad Campaigns?

### v1

@@ 4.2 Efficient Influence Function
Part (i) of Theorem 4.3 states that $\widehat{\theta}_{\textnormal{EIF,a}}(\gamma_{a})$ is conditionally doubly robust in that if $\widehat{\rho}_{a}^{(k)}({\bf v})$ is consistent, $\widehat{\theta}_{\textnormal{EIF,a}}(\gamma_{a})$ is consistent when either $\widehat{\pi}^{(k)}({\bf x})$ or $\widehat{\mu}_{a}^{(k)}({\bf x})$, but not necessarily both, is consistent. Part (ii) states that if all the nuisance parameters are estimated consistently at the rates in equations (4.4a)-(4.4c), $\widehat{\theta}_{\textnormal{EIF,a}}(\gamma_{a})$ is asymptotically normal and semiparametrically efficient. We remark that when transportability holds, our result recovers Theorem 5 of Zeng et al., (2023), which does not require equation (4.4c). In other words, equation (4.4c) can be viewed as the cost of violating transportability. Finally, for inference, parts (ii) and (iii) of Theorem 4.3 imply an asymptotically valid, $1-\alpha$ CI of $\theta_{a}(\gamma_{a})$ is $\widehat{\theta}_{\textnormal{EIF,a}}(\gamma_{a})\pm z_{1-\alpha/2}\sqrt{\widehat{\sigma}_{\textnormal{EIF,a}}^{2}(\gamma_{a})}$ where $z_{1-\alpha/2}$ is the $21-\alpha/$ 2 quantile of the standard normal distribution.
@@
Equation (4.4c) requires that we not only consistently estimate the outcome regression $\rho_{a}({\bf v})$, but also estimate it at a sufficiently fast rate. If $\rho_{a}({\bf v})$ is a finite-dimensional, parametric function as in Theorem 4.1, equation (4.4c) is satisfied with a parametric estimator. However, if $\rho_{a}({\bf v})$ is a non-parametric function, for instance a Lipschitz function with smoothness $m$, we effectively need $m\geq d/2$ where $d$ is the dimension of ${\bf V}_{i}$ to satisfy (4.4c); see Kennedy, (2023). In particular, we cannot estimate $\rho_{a}({\bf v})$ at a slow rate in hopes that another estimator of the nuisance function can “compensate” for the slow rate; this is referred to as the mixed bias property or rate double robustness (e.g., Rotnitzky et al., (2020), Kennedy, (2022)). In contrast, a common approach to satisfy equation (4.4a) is to obtain data from an RCT where the propensity score is known a priori and to estimate the outcome regression using a flexible, machine learning method, which may converge slowly.
5 Calibrating the Sensitivity Parameters
5.1 A Three-Step Calibration Procedure
Suppose that an investigator estimates the TATE for a range of $(\gamma_{0},\gamma_{1})$ using the methods in Sections 4.1 or 4.2 and finds that some estimates are statistically significant at the $\alpha=0.05$ level while others are not. Perhaps, the investigator went a step further and found the “smallest” $(\gamma_{0},\gamma_{1})$ where the estimate becomes insignificant (or significant). In both cases, a natural question to ask is what values of the sensitivity parameters indicate plausible deviations between the source and target populations. More generally, finding a plausible, “reference” set of sensitivity parameters has been a long-standing question in sensitivity analysis and this task is often referred to as calibration or benchmarking (Cinelli and Hazlett,, 2020; Huang, 2024b, ). In this section, we present a calibration procedure that generates “reference” values of the sensitivity parameters using an idea underlying design sensitivity (Rosenbaum,, 2004, 2020); see Section 5.2 for more discussion. The calibration procedure is roughly divided into three steps, and a visualization of this procedure for Philadelphia county is in Figure D.1 of the Appendix.
The first step partitions the source data from Aggarwal et al., (2023) into “blue collar” states (PA, MI, WI), denoted as ${\cal I}_{s_{1}}$, and non-blue collar states (AZ, NC), denoted as ${\cal I}_{s_{2}}$. In general, the two partitions ${\cal I}_{s_{1}}$ and ${\cal I}_{s_{2}}$ of the source data are chosen to have unmeasurable differences between them. For example, with the blue collar and non-blue collar states, there may be differences in socioeconomic status, labor markets, and region-specific politics (i.e., Midwestern states versus non-Midwestern states) not captured by ${\bf V}_{i}$.
@@ 5.1 A Three-Step Calibration Procedure
The second step temporarily treats the voters in non-blue collar states as the “proxy” target population and constructs two $1-\alpha$ CIs of the TATE in this proxy target population:
@@
  • (Our Transfer Learning Approach): We treat the blue collar states as the “proxy” source population and use the methods in Sections 4.1 or 4.2 to infer the proxy target population. We denote the resulting confidence interval as ${\widehat{\text{CI}}}_{s_{1}\to s_{2}}(\gamma_{0},\gamma_{1};1-\alpha)$.
@@
  • (Standard Approach): Using the proxy target data only (i.e., data from non-blue collar states), we compute a valid $(1-\alpha)$ CI of the TATE, say the Wald confidence interval based on the difference-in-means estimator, and denote it as $\widehat{{\text{CI}}}_{s_{2}}(1-\alpha)$.
A key insight is that if Assumption 2.1 holds, the CI from the standard approach (i.e., $\widehat{{\text{CI}}}_{s_{2}}(1-\alpha)$) will cover the ATE in the proxy target population. In our example where the source data is generated from a randomized experiment, Assumption 2.1 holds by design and $\widehat{{\text{CI}}}_{s_{2}}(1-\alpha)$ has valid coverage of the ATE among the non-blue collar states. In contrast, the CI from transfer learning ${\widehat{\text{CI}}}_{s_{1}\to s_{2}}(\gamma_{0},\gamma_{1};1-\alpha)$ may not cover the true ATE in the proxy target population unless $(\gamma_{0},\gamma_{1})$ are chosen correctly to measure the true difference between the proxy source and the proxy target population.
Recognizing that $\widehat{{\text{CI}}}_{s_{2}}(1-\alpha)$ is a valid CI with only Assumption 2.1, we “calibrate” the transported ${\widehat{\text{CI}}}_{s_{1}\to s_{2}}(\gamma_{0},\gamma_{1};1-\alpha)$ by keeping the values of $(\gamma_{0},\gamma_{1})$ where both CIs overlap, or formally $\mathcal{C}_{1}=\{(\gamma_{0},\gamma_{1})\mid\widehat{{\text{CI}}}_{s_{1}\to s_{2}}(\gamma_{0},\gamma_{1};1-\alpha)\cap\widehat{{\text{CI}}}_{s_{2}}(1-\alpha)\neq\emptyset\}$. In other words, $\mathcal{C}_{1}$ represent plausible sensitivity parameters that capture the true unmeasured differences between the two populations. Also, conducting the original sensitivity analysis of the TATE with the sensitivity parameters in $\mathcal{C}_{1}$ restricts the original analysis to unobservable differences that the investigator designed by creating meaningful/explainable partitions of the source population.
We conclude the section with some brief remarks on implementing the calibration procedure. First, we recommend repeating the three steps above, but with the roles of the proxy target and the source populations reversed, yielding another set $\mathcal{C}_{2}$ of sensitivity parameters. The intersection $\mathcal{C}=\mathcal{C}_{1}\cap{\cal C}_{2}$ consists of what we call the plausible values of the sensitivity parameters. Also, if any value in $\mathcal{C}$ leads to statistically significant effects, we say that the treatment effect is sensitive for the effect. In contrast, if none of the values in $\mathcal{C}$ lead to statistically significant effects, we say that the treatment effect is insensitive for the effect. Second, to better account for sampling variability between the original target and source populations, we recommend re-scaling the standard errors in the calibration procedure to match the ratio of the size of the original source population and the original target population; see Section D of the Appendix for details.
5.2 Comparison to Existing Methods
The inspiration for the calibration procedure comes from Rosenbaum’s extensive works on design sensitivity, specifically the notion of a “favorable situation” in Chapter 15 of Rosenbaum, (2020) to benchmark designs of observational studies in terms of robustness against unmeasured confounding. Specifically, the proxy target population in our calibration procedure is similar to Rosenbaum’s favorable situation in that the investigator chooses a situation where the ATE is known and is free from hidden bias. In our setup, the proxy target’s ATE is not unknown, but can be unbiasedly estimated under Assumption 2.1 only; we do not need transportability, i.e., the assumption that we are conducting sensitivity analysis for, to identify the proxy target’s ATE. Also, roughly speaking, the set $\mathcal{C}_{1}$ contains sensitivity parameters that will “reject” towards the favorable situation. However, unlike design sensitivity, which, in our setting, would be roughly equivalent to taking the “limit” of the set $\mathcal{C}_{1}$ when sample size increases and using this limit to benchmark designs of observational studies, we directly use $\mathcal{C}_{1}$ to calibrate/benchmark our sensitivity parameters.
Another common way of calibrating sensitivity parameters is to omit an observed covariate (e.g., (Hsu and Small,, 2013; Cinelli and Hazlett,, 2020; Ek and Zachariah,, 2023; Huang, 2024b, )). But, as discussed in Section 6.2 of Cinelli and Hazlett, (2020), this can lead to a misleading understanding of the magnitude of unmeasured confounding. Our calibration procedure avoids this issue by using the same covariates for sensitivity analysis and calibration.
Finally, similar to other calibration procedures, our calibration procedure cannot “…rule out the existence of killer confounders” (Huang, 2024b, ) and, if used incorrectly, can give misleading conclusions about the magnitude of the violation of transportability. For example, we believe that an ineffective use of our procedure is to create a partition where there are no unmeasurable differences between the two proxy populations. Nevertheless, compared to omitting a confounder, we believe creating dissimilar partitions of the study sample can be a promising way to calibrate unobservable differences between the source and the target populations.

### v2

@@ 4.2 Efficient influence function
Part (i) of Theorem 4.3 states that $\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a})$ is conditionally doubly robust in that if $\widehat{\rho}_{a}^{(k)}$ is consistent, $\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a})$ is consistent when either $\widehat{\pi}^{(k)}({\bf x})$ or $\widehat{\mu}_{a}^{(k)}$, but not necessarily both, is consistent. Part (ii) states that if all the nuisance functions are estimated consistently at the rates in equations (4.4a)-(4.4c), $\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a})$ is asymptotically normal and semiparametrically efficient. We remark that when transportability holds, our result recovers Theorem 5 of Zeng et al., (2023), which does not require equation (4.4c); see below for more discussions. Finally, Theorem 4.3 implies that an asymptotically valid, $1-\alpha$ CI of $\theta_{a}(\Gamma_{a})$ is $\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a})\pm z_{1-\alpha/2}\sqrt{\widehat{\sigma}_{\textnormal{EIF,a}}^{2}(\Gamma_{a})}$ where $z_{1-\alpha/2}$ is the $21-\alpha/$ 2 quantile of the standard normal distribution.
@@
Equation (4.4c) requires that we not only consistently estimate the outcome regression $\rho_{a}({\bf v})$, but also estimate it at a sufficiently fast rate. If the true $\rho_{a}({\bf v})$ is a parametric function as in Theorem 4.1, equation (4.4c) is satisfied with a parametric estimator. Otherwise, we cannot estimate $\rho_{a}({\bf v})$ at a slow, nonparametric rate in hopes that another estimator of the nuisance function can “compensate” for the slow rate; this is referred to as the mixed bias property or rate double robustness (e.g., Rotnitzky et al., (2020), Kennedy, (2022)). In contrast, one approach to satisfy equation (4.4a) is to obtain data from an RCT where the propensity score is known a priori and estimate the outcome regression using a supervised machine learning method, which may converge slowly. More broadly, equation (4.4c) can be viewed as the cost of violating transportability. Intuitively, we incur this cost because the sensitivity model (3.1) is based on shifting the outcome distribution and if $\rho_{a}$ is poorly estimated, the sensitivity model is also incorrectly specified, which ultimately leads to a poor estimate of the TATE; see Section C of the Appendix for more details.
4.3 Estimation of nuisance functions
This section briefly discusses estimation of the nuisance parameters, specifically $\rho_{a}$ and $\omega$. For the other, “classical” nuisance functions (i.e., propensity score $\pi$ and the outcome regression function $\mu_{a}$), we echo the modern recommendation of using the investigator’s favorite classification and regression models. Note that if the source data is from an RCT, investigators should use the design of the RCT to estimate $\pi$.
The regression function $\rho_{a}$ can be estimated in a couple of different ways and we highlight each approach through the equalities below:
$\rho_{a}({\bf v})$ $=\mathbb{E}\{\mu_{a}({\bf X}_{i})\mid{\bf V}_{i}={\bf v},S_{i}=1\}$ (4.5)
$=\mathbb{E}\left[\left\{\dfrac{A_{i}\mathds{1}(A_{i}=a)}{\pi({\bf X}_{i})}+\dfrac{(1-A_{i})\mathds{1}(A_{i}=1-a)}{1-\pi({\bf X}_{i})}\right\}Y_{i}\mid{\bf V}_{i}={\bf v},S_{i}=1\right]$ (4.6)
$=\mathbb{E}\left[\left\{\dfrac{A_{i}\mathds{1}(A_{i}=a)}{\pi({\bf X}_{i})}+\dfrac{(1-A_{i})\mathds{1}(A_{i}=1-a)}{1-\pi({\bf X}_{i})}\right\}\{Y_{i}-\mu_{a}({\bf X}_{i})\}+\mu_{a}({\bf X}_{i})\mid{\bf V}_{i}={\bf v},S_{i}=1\right].$ (4.7)
The first equality (4.5) suggests estimating $\rho_{a}$ by regressing the predicted outcome $\widehat{\mu}_{a}({\bf X}_{i})$ on ${\bf V}_{i}$. The second equality (4.6) suggests regressing an inverse-probability-weighted (IPW) outcome $\left[A_{i}\mathds{1}(A_{i}=a)/\widehat{\pi}({\bf X}_{i})+(1-A_{i})\mathds{1}(A_{i}=1-a)/\{1-\widehat{\pi}({\bf X}_{i})\}\right]Y_{i}$ on ${\bf V}_{i}$. The third equality (4.7) suggests regressing an augmented IPW outcome,
$\left[A_{i}\mathds{1}(A_{i}=a)/\widehat{\pi}({\bf X}_{i})+(1-A_{i})\mathds{1}(A_{i}=1-a)/\{1-\widehat{\pi}({\bf X}_{i})\}\right]\{Y_{i}-\widehat{\mu}_{a}({\bf X}_{i})\}+\widehat{\mu}_{a}({\bf X}_{i})$
on ${\bf V}_{i}$. Under the first and the third approaches, the rate of convergence of $\widehat{\rho}_{a}$ is dependent of the rate of $\widehat{\mu}_{a}$ (Kennedy,, 2023). Under the second approach, the rate of convergence of $\widehat{\rho}_{a}$ is independent of the rate of convergence of $\widehat{\mu}_{a}$. We remark that when all covariates are discrete and $\pi$ and $\mu_{a}$ are estimated by taking means within subgroups defined by the covariates, the three approaches are equivalent.
For estimating the density ratio $w({\bf v})$, we recommend entropy balancing methods (Hainmueller,, 2012; Josey et al.,, 2022; Chen et al.,, 2023), which obtains $\widehat{w}({\bf V}_{i})$ as solutions to the following constrained optimization problem,
$\begin{split}&\underset{w_{i}}{\rm argmin}\sum_{i\in{\cal I}_{s}}w_{i}{\textnormal{log}}(w_{i}),\quad{}\textnormal{s.t. }\dfrac{1}{n_{s}}\sum_{i\in{\cal I}_{s}}w_{i}{\bf V}_{i}=\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}{\bf V}_{i}.\end{split}$ (4.8)
If the true $\mathbb{P}(S_{i}=1\mid{\bf V}_{i})$ is a logistic regression model and the parameters of the model are identified, the probability limit of the weights in (4.8) is equal to $\mathbb{P}(S_{i}=0\mid{\bf V}_{i})\mathbb{P}(S_{i}=1)/\{\mathbb{P}(S_{i}=1\mid{\bf V}_{i})\mathbb{P}(S_{i}=0)\}$. Otherwise, the weights in (4.8) generally have favorable, finite-sample properties (e.g., Chen et al., (2023)). For more discussions on estimating $w({\bf v})$, see Section C of the Appendix.
5 Calibrating sensitivity parameters
5.1 Definition of a sensitive effect and motivation for calibration
Section 4 provides two procedures to estimate the TATE for a given value of $(\Gamma_{0},\Gamma_{1})$ and allows investigators to study the change of the TATE as it moves away from $(\Gamma_{0},\Gamma_{1})=(1,1)$, i.e., the setting where transportability holds. Traditionally, investigators consider several $(\Gamma_{0},\Gamma_{1})$s that are not equal to $(1,1)$ and assess whether the statistical conclusion about the TATE changes between $(\Gamma_{0},\Gamma_{1})=(1,1)$ and other $(\Gamma_{0},\Gamma_{1})$s. A bit more formally, let ${\cal C}\subset\mathbb{R}^{+}\times\mathbb{R}^{+}\neq\{(1,1)\}$ denote the set of $(\Gamma_{0},\Gamma_{1})$s that the investigator is considering for the sensitivity analysis. Following the literature, we say the TATE is sensitive to transportability if the decision to reject the null hypothesis of no effect at the significance level $\alpha$ changed between $(\Gamma_{0},\Gamma_{1})=(1,1)$ and another value of $(\Gamma_{0},\Gamma_{1})\in{\cal C}$.
Definition 5.1 (Sensitivity to Transportability)
Consider a significance level $\alpha$ and the set ${\cal C}\subset\mathbb{R}^{+}\times\mathbb{R}^{+}\neq\{(1,1)\}$. For a given $(\Gamma_{0},\Gamma_{1})$, let $\widehat{\textnormal{CI}}(\Gamma_{0},\Gamma_{1};1-\alpha)$ denote a $1-\alpha$ CI of the TATE from Section 4. The TATE is sensitive to transportability in ${\cal C}$ if there exists $(\Gamma_{0},\Gamma_{1})\in{\cal C}$ such that either of the following holds:
$\begin{split}\text{(i) from significant to insignificant:}\quad{}&0\not\in\widehat{\textnormal{CI}}(1,1;1-\alpha)\text{ and }0\in\widehat{\textnormal{CI}}(\Gamma_{0},\Gamma_{1};1-\alpha);\\ \text{(ii) from insignificant to significant:}\quad{}&0\in\widehat{\textnormal{CI}}(1,1;1-\alpha)\text{ and }0\not\in\widehat{\textnormal{CI}}(\Gamma_{0},\Gamma_{1};1-\alpha).\end{split}$
If neither (i) nor (ii) holds, the TATE is insensitive to transportability in ${\cal C}$.
Some investigators have a well-defined ${\cal C}$ based on their belief about the unmeasured difference between the source and the target populations in the odds ratio scale. But in general, specifying a reasonable, “reference” set of sensitivity parameters has been a long-standing question in the literature on sensitivity analysis and this task is often referred to as calibration or benchmarking (Cinelli and Hazlett,, 2020; Huang, 2024b, ). One popular approach to generate the reference values is to omit an observed covariate (e.g., Hsu and Small, (2013); Cinelli and Hazlett, (2020); Ek and Zachariah, (2023); Huang, 2024b ) and conduct the sensitivity analysis with the values of the sensitivity parameters that are comparable to the effects of the omitted covariate on the outcome or the treatment. But, as discussed in Section 6.2 of Cinelli and Hazlett, (2020), this can lead to a misleading understanding of the magnitude of unmeasured confounding, especially when the omitted variable is correlated with other confounders.
In this section, we present a data-driven calibration procedure that generates a reference, calibrated set of the sensitivity parameters using an idea from design sensitivity (Rosenbaum,, 2004, 2020). Briefly, design sensitivity is used to benchmark designs of observational studies in terms of robustness against unmeasured confounding by measuring the limiting power to accept a particular type of alternative hypothesis, referred to as a “favorable situation” (Chapter 15 of Rosenbaum, (2020)). While Rosenbaum created the favorable situation from parametric models, we create it using the source data. Also, our calibrated set is a finite-sample, two-dimensional analog of Rosenbaum’s design sensitivity parameter in that the sensitivity parameters in the calibrated set lead to “accepting” the favorable situation created from data. Importantly, our calibration procedure avoids the issue from the omitted variables approach above by using the same covariates for both sensitivity analysis and calibration.
We state the calibration procedure in Section 5.2 and describe the rationale in Section 5.3.
5.2 Calibration procedure
@@
The calibration procedure is divided into three steps. The first step partitions the source data from Aggarwal et al., (2023) into the rust belt states (PA, MI, WI), denoted as ${\cal I}_{s_{1}}$, and the sun belt states (AZ, NC), denoted as ${\cal I}_{s_{2}}$. The second step temporarily treats the voters in the sun belt states as the “proxy” target population and constructs two $1-\alpha$ CIs of the TATE for it:
@@
  • (Our Transfer Learning Approach): We treat the rust belt states as the “proxy” source population and use the methods in Section 4 to infer the TATE in the sun belt states (i.e., the proxy target population). We denote the resulting confidence interval as ${\widehat{\text{CI}}}_{s_{1}\to s_{2}}(\Gamma_{0},\Gamma_{1};1-\alpha)$.
@@
  • (The Standard Approach): Using the data from the sun belt states only (i.e., the proxy target population), we compute a valid $1-\alpha$ CI of the TATE, say the Wald confidence interval based on the difference-in-means estimator, and denote it as $\widehat{{\text{CI}}}_{s_{2}}(1-\alpha)$.
The third step keeps the values of $(\Gamma_{0},\Gamma_{1})$ where the CIs from both approaches overlap, or formally, $\mathcal{C}_{1}=\{(\Gamma_{0},\Gamma_{1})\mid\widehat{{\text{CI}}}_{s_{1}\to s_{2}}(\Gamma_{0},\Gamma_{1};1-\alpha)\cap\widehat{{\text{CI}}}_{s_{2}}(1-\alpha)\neq\emptyset\}$. We repeat the three steps above, but with the roles of the proxy target and proxy source populations reversed, yielding another set of sensitivity parameters ${\cal C}_{2}$. The intersection of the two sets, ${\cal C}={\cal C}_{1}\cap{\cal C}_{2}$, is the data-driven, calibration set of sensitivity parameters. Further computational details and a step-by-step algorithm are provided in Section D of the Appendix.
5.3 The rationale behind the calibration procedure
The first step partitions the source data into two subsets ${\cal I}_{s_{1}}$ and ${\cal I}_{s_{2}}$ such that there are scientifically meaningful, unobserved differences between them. For example, there are meaningful differences in socioeconomic status, labor markets, and region-specific politics between the sun belt states and the rust belt states and these differences are not measured by ${\bf V}_{i}$. We remark that investigators can choose other partitions that are interpretable; see below for more discussion.
After partitioning the data, the next two steps find $(\Gamma_{0},\Gamma_{1})$s that quantify the unmeasured differences between ${\cal I}_{s_{1}}$ and ${\cal I}_{s_{2}}$. This is accomplished by finding a set of $(\Gamma_{0},\Gamma_{1})$s such that the transported ATE (i.e., “Our Transfer Learning Approach” above) is close to the true ATE of the proxy target population, up to sampling error. Note that the true TATE in the proxy target population can be inferred with only Assumption 2.1, specifically using the $1-\alpha$ CI from the standard approach above. Then, the resulting set ${\cal C}$ represents the magnitude of the unmeasured differences between the two subsets ${\cal I}_{s_{1}}$ and ${\cal I}_{s_{2}}$ as the values in ${\cal C}$ correctly transported from the proxy source data to match the true ATE of the proxy target population.
It’s important to recognize that the set ${\cal C}$ obtained from the calibration procedure is not the true unmeasurable differences between the original source population and the original target population, for instance the unmeasured differences between the 2020 and 2024 elections. Similarly, using the calibrated set ${\cal C}$ in the sensitivity analysis does not imply that the true unmeasured differences between the original source and the target populations can be estimated from the unmeasured differences between the two subsets of the original source data. As mentioned in Section 3, the $(\Gamma_{0},\Gamma_{1})$ that parametrizes the unmeasured differences between the original source and the target populations cannot be identified or estimated.
Instead, akin to the usual approach of conducting sensitivity analysis based on a set ${\cal C}$ informed by the investigator’s prior belief, the calibrated set $\mathcal{C}$ is a data-driven approach to generate another interpretable set of sensitivity parameters. For example, if the TATE of the 2024 election is sensitive with respect to the set $\mathcal{C}$ generated from the investigator’s prior belief, it suggests that the unmeasured differences that are as large as those hypothesized by the investigator can overturn the conclusion about the TATE in the 2024 election. Similarly, if the TATE of the 2024 election is sensitive with respect to the calibrated set $\mathcal{C}$, it suggests that the unmeasured differences that are as large as those between the sun belt states and the rust best states in the 2020 election can overturn the conclusion about the TATE in the 2024 election. In short, our calibration procedure is another approach to understand and interpret the sensitivity parameters that is based on the observed data.
We also briefly mention a subtle point about the sample size and sampling uncertainty in the calibration procedure. Technically speaking, the partitioning step has a different sample size than that for the original analysis, which leads to different magnitudes of sampling uncertainty. Section D of the Appendix discusses how we re-scale the standard errors and conduct downsampling in the calibration procedure so that the sampling uncertainty is comparable between the calibration procedure and the original analysis. A broader discussion about sampling uncertainty in transfer learning is in Section 7.
Finally, as mentioned in the beginning of this section, investigators can choose other partitions of the source data in the first step. But, some partitions are more useful than others. For example, a random partition of the source data such that there are no unmeasurable differences between the two subsets is not meaningful. Nevertheless, between two non-random partitions, some investigators may find one partition to be more interpretable than the other. In fact, the investigators’ unrestricted ability to choose a partition is a useful feature of our calibration procedure compared to the omitted variable approach where the investigators are restricted to the list of observed confounders or usually focus on the “strongest” confounder. In general, compared to the calibration procedure based on omitting a covariate, we believe creating dissimilar partitions of the source is a promising way to study unobservable differences between the source and the target populations.